### v2

# Quantifying the Balinski-Young Theorem: Structure and Probability of Quota Violations in Divisor Methods for Three States

Abstract.
The apportionment problem asks how to assign representation to states based on their populations. That is, given census data and a fixed number of seats, how many seats should each state be assigned? Various algorithms exist to solve the apportionment problem, but by the Balinski-Young Theorem, every such algorithm will be flawed in some way. This paper focuses on divisor methods of apportionment, where the possible flaws are known as quota violations. This paper presents a detailed analysis of quota violations that can arise under divisor methods for three states. The study focuses on quota violations in Adams’s, Jefferson’s, Dean’s, and the Huntington-Hill methods when allocating $M$ seats. Theoretical results are proved about the behavior of these methods, particularly focusing on the types of quota violations that may occur, their frequency, and their structure. The paper then introduces tests to detect quota violations, which are then employed to construct a probability function which calculates the likelihood of such violations occurring given an initial three state population vector whose components follow varying distributions.
1. Introduction
The apportionment problem - dividing a fixed number of indivisible resources among states in proportion to their populations - lies at the intersection of political science, combinatorics, and probability. A key class of apportionment methods are divisor methods, which are a group of simple methods that avoid unfair results known as apportionment paradoxes [BY01]. Divisor methods have been widely employed in the United States for assigning seats in House of Representatives, with Jefferson’s, Webster’s method, or the Huntington-Hill methods used for most of the country’s history. The Huntington-Hill method is currently used to apportion the House of Representatives.
While divisor methods avoid paradoxes, these methods do not always satisfy the quota rule, which requires that each state receive an allocation equal to either the floor or ceiling of its exact proportional share. This phenomenon of imperfect apportionment is in fact guaranteed by the Balinski-Young Theorem [BY01], which establishes that no apportionment method can simultaneously satisfy the quota rule and avoid paradoxes.
Since divisor methods can violate quota, and as flawed methods are unavoidable, a good question to then ask is how often will a quota violation happen, and can quota violations be predicted in advance given states’ populations? This paper answers that question for three states. Specifically, this paper investigates the frequency and structure of quota violations under divisor methods in three states in the methods of Adams’s, Jefferson’s, Webster’s, Dean’s, and Huntington-Hill. The paper characterizes how quota violations occur and provides formulas to calculate the probability that a divisor method violates quota.
2. Preliminary Notation and Definitions
Definition 2.1.
Given a strictly increasing divisor function $d:\mathbb{N}\to\mathbb{R}_{\geq 0}$, populations $(p_{1},\dots,p_{n})$ and a fixed number of seats $M$, an apportionment function $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$, is then defined as follows:
  1. 1.
    Assign each state zero seats.
  2. 2.
    For each state calculate their priority value, $\frac{p_{i}}{d(r_{i})}$, where $r_{i}$ is the number of seats currently assigned to state $i$.
  3. 3.
    Assign to the state with the highest priority value another seat.
  4. 4.
    Repeat steps (2.) and (3.) until $M$ seats have been apportioned.
If $d(0)=0$, use the convention that division by zero results in infinity, and as such all states immediately receive one seat. Equivalently, if $d(0)=0$, replace the first step with assigning each state one seat rather than zero.
The most common and arguably reasonable divisor methods are the “five workable methods” [BY77, BY01, Hunt21] defined by the following divisor functions:
Name Divisor Function
Adams $d(s)=s$
Jefferson $d(s)=s+1$
Webster $d(s)=s+0.5$
Huntington-Hill $d(s)=\sqrt{s(s+1)}$
Dean $d(s)=\dfrac{2s(s+1)}{2s+1}$
Important values are the standard quotas and standard divisors. Let $P=\sum_{i=1}^{n}p_{i}$ be the total population. Then, the standard divisor is defined as $SD=\frac{P}{M}$, and state $i$’s standard quota is $q_{i}=\frac{p_{i}}{SD}.$ The standard quota is a percentage $\frac{p_{i}}{P}$ of $M$ and is a state’s theoretical fair share of the number of seats. As such each state should be apportioned roughly their standard quota. Since the standard quota is rarely an integer, states should be apportioned either their upper quota, $\lceil q_{i}\rceil$, or their lower quota, $\lfloor q_{i}\rfloor$.
A quota violation occurs whenever a state is assigned more seats than their upper quota or fewer seats then their lower quota. Specifically, a lower quota violation occurs when $a_{i}<\lfloor q_{i}\rfloor$ and an upper quota violation occurs when $a_{i}>\lceil q_{i}\rceil$. Informally, a quota violation is when a state is apportioned too many or too few seats.
The Balinski-Young Theorem [BY01] guarantees that all apportionment methods either have quota violations or have strange, unfair occurrences known as apportionment paradoxes. Divisor methods have quota violations but avouid paradoxes [BY01].
3. Results
The main results of the paper are:
Theorem 5.1.
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
Theorem 5.5.
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Adams’s, Dean’s, or the Huntington Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
Theorem 5.6.
(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. (1)
    $q_{3}d(\lfloor q_{1}\rfloor)<q_{1}d(\lfloor q_{3}\rfloor-1)$
  2. (2)
    $q_{3}d(\lfloor q_{2}\rfloor)<q_{2}d(\lfloor q_{3}\rfloor-1)$
  3. (3)
    $\lceil q_{1}\rceil+\lceil q_{2}\rceil+\lfloor q_{3}\rfloor=M+1$ (equivalently for $q_{1},q_{2}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1)$.
Theorem 5.9.
(Probability of Lower Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and for apportionment method $A$ equal to Adams, Dean, or Huntington-Hill, the probability of a lower quota violation is given by:
$\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
where $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ denotes the feasible region of the Lower Quota Violation Criteria Test (5.6) defined in Theorem 5.7.
Theorem 6.4.
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
Theorem 6.5.
(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. (1)
    $q_{3}d(\lfloor q_{1}\rfloor)>q_{1}d(\lceil q_{3}\rceil)$
  2. (2)
    $q_{3}d(\lfloor q_{2}\rfloor)>q_{2}d(\lceil q_{3}\rceil)$
  3. (3)
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
Theorem 6.7.
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
$\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
where $\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ denotes the feasible region of the Upper Quota Violation Criteria Test (6.5) defined in Theorem 6.6.
4. Background Information
The following are simple lemmas, known results, and common assumptions that are utilized in this paper. Throughout the paper, let $M$ be the number of seats and $n$ the number of states. Let $A(p_{1},\ldots,p_{n})$ be the divisor method apportionment for state populations $p_{i}$, with $A$ an arbitrary divisor method.
4.1. Assumptions
Assume there are no ties in priority values; in particular this includes assuming no ties in states’ populations. Additionally, to guarantee at least one seat per state assume $M\geq n$.
4.2. Introductory Lemmas and Definitions
Lemma 4.1.
[…]
Lemma 4.2.
(Homogeneity) A divisor method is preserved under proportionality. If $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$ then for any scalar, $\lambda\in\mathbb{R}_{>0}$, $A(\lambda p_{1},\dots,\lambda p_{n})=(a_{1},\dots,a_{n})$. In particular, apportioning the standard quotas is the same as apportioning the populations: $A(p_{1},\dots,p_{n})=A(q_{1},\dots,q_{n})$. [BY01].
Definition/Lemma 4.3.
Define a divisor apportionment method $\tilde{A}(p_{1},\ldots,p_{n})$ on $n$ states and $M$ seats by finding a modified divisor $D\in\mathbb{R}_{>0}$ such that $D$ causes $\sum_{i=0}^{n}a_{i}=M$, for $a_{i}$ calculated as follows:
  1. (1)
    For each state, calculate a modified quota $m_{i}=\frac{p_{i}}{D}.$
  2. (2)
    For each state, assign seats by modified-rounding: round up if $m_{i}\geq d(\lfloor m_{i}\rfloor)$; round down if less.
The equivalence of the apportionment methods $A(p_{1},\ldots,p_{n})$ in Definition 2.1 and $\tilde{A}(p_{1},\ldots,p_{n})$ is classical. See [BY01], [Hunt21], and [Pukelsheim2017].
Theorem 4.4.
No divisor method will have both upper and lower quota violations for any $M$. [BY01, KS18].
Theorem 4.5.
(No Violation Theorem) From [BY01], a summary of cases when no quota violations occur is as follows:
  • (a)
    For $n=2$ states, no upper or lower quota violations occur in any divisor methods.
  • (b)
    For $n=3$ states, Webster’s Method is the unique method such that upper or lower quota violations do not occur.
  • (c)
    Adams’s Method is the unique method where upper quota violations do not occur for all $n$.
  • (d)
    Jefferson’s Method is the unique method where lower quota violations do not occur for all $n$.
5. Structure and Probability of Quota Violations for Three States in Adams’s, Dean’s, and Huntington-Hill’s Method
We start by providing a detailed study of lower quota violations in Adams’s, Dean’s, and Huntington-Hill’s method when apportioning three states. Webster’s and Jefferson’s method have no lower quota violations by Theorem 4.5.
5.1. Characterization of Quota Violations in Three States
By the following theorem, only lower quota violations can occur in Adams’s, Dean’s, and Huntington-Hill’s method for three states.
Theorem 5.1.
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
Proof.
Let $p_{i}$ and $q_{i}$ be state $i$’s populations and standard quotas respectively. Then $A(p_{1},p_{2},p_{3})=A(q_{1},q_{2},q_{3})=(a_{1},a_{2},a_{3})$. Let $d(s)\in[s,s+\frac{1}{2}]$ for all $s$. Assume for contradiction there exists an upper quota violation and, without loss of generality, the violation is on state one. Then, $a_{1}\geq\lceil q_{1}\rceil+1$, and if $q_{1}$ is not an integer $a_{1}\geq\lfloor q_{1}\rfloor+2$.
Write each $q_{i}$ as $\lfloor q_{i}\rfloor+d_{i}$, where $d_{i}$ is the decimal part. Since an upper quota violation exists by assumption, by Theorem 4.4, no lower quota violation exists. Therefore $a_{2}\geq\lfloor q_{2}\rfloor$ and $a_{3}\geq\lfloor q_{3}\rfloor$. Then if $q_{1}$ is not an integer:
$\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor+d_{1}+d_{2}+d_{3}$ $=q_{1}+q_{2}+q_{3}$
$=M$
$=a_{1}+a_{2}+a_{3}$
$\geq\lceil q_{1}\rceil+1+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor$
$=\lfloor q_{1}\rfloor+2+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor$
Which implies $\sum d_{i}\geq 2$. Since $d_{i}\in[0,1)$, at least two $d_{i}$ are greater than $\frac{1}{2}$. Since $d(s)\in[s,s+\frac{1}{2}]$, there are at least two roundups when assigning apportionment for the modified divisor $D=SD$, resulting in two states assigned their upper quota and one state with an unknown assignment of either its lower or upper quota.
Similarly, if $q_{1}$ is an integer, then:
$\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor+d_{2}+d_{3}$ $=q_{1}+q_{2}+q_{3}$
$=a_{1}+a_{2}+a_{3}$
$\geq\lceil q_{1}\rceil+1+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor$
$=\lfloor q_{1}\rfloor+1+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor$
Which implies $d_{2}+d_{3}\geq 1$ and therefore at least one of $d_{2}$, $d_{3}$ is greater than $\frac{1}{2}$. This results in one of those two states apportioned its upper quota and state one apportioned its upper quota. The final apportionment is then two states assigned their upper quota and one state with an unknown assignment of either its lower or upper quota.
As such, for $q_{1}$ an integer or not, two states are assigned their upper quota and one state has an unknown assignment of either their lower or upper quota. If this third state is assigned its lower quota, pick $d=SD$ and there is no upper quota violation. If instead the upper quota is assigned, pick $D>SD$. This cannot create an upper quota violation as increasing $D$ reduces apportionments. Both cases contradict the assumption that an upper quota violation occurred. ∎
The proof above allows for lower quota violations to occur in $n=3$ states and indeed they do. The following three lemmas completely classify the lower quota violations in Adams’s, Dean’s and Huntington-Hill’s apportionment methods.
Lemma 5.2.
In a divisor apportionment method with $M$ seats, $n=3$ states, and standard quotas $q_{i}$, if a lower quota violation occurs then the apportionment of the state with the violation is $\lfloor q_{i}\rfloor-1$ and the other two states are assigned their upper quotas.
Proof.
Assume without loss of generality that the lower quota violation occurs on state 3. The claim is that the final apportionment is then $a_{1}=\lceil q_{1}\rceil$, $a_{2}=\lceil q_{2}\rceil$, and $a_{3}=\lfloor q_{3}\rfloor-1$. Since the first two states cannot be apportioned more than the ceiling of their quota by Theorem 4.4, $a_{1}\leq\lceil q_{1}\rceil$ and $a_{2}\leq\lceil q_{2}\rceil$. Similarly state 3 cannot have an apportionment larger than $\lfloor q_{3}\rfloor-1$ by the existence of a lower quota violation, so $a_{3}\leq\lfloor q_{3}\rfloor-1$.
Write $q_{i}=\lfloor q_{i}\rfloor+d_{i}$, where $d_{i}$ is the decimal part. Then $M=\sum q_{i}=\sum\lfloor q_{i}\rfloor+d_{i}=\sum\lfloor q_{i}\rfloor+\sum d_{i}$. In particular, $\sum d_{i}=M-\sum\lfloor q_{i}\rfloor$ is an integer and since each $d_{i}\in[0,1),\sum d_{i}\in\{0,1,2\}.$
Then
$a_{1}+a_{2}=M-a_{3}\geq M-(\lfloor q_{3}\rfloor-1)=\sum\lfloor q_{i}\rfloor+\sum d_{i}-\lfloor q_{3}\rfloor+1=\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\sum d_{i}+1$
Proceed for each case of $\sum d_{i}\in\{0,1,2\}:$
If $\sum d_{i}=2,a_{1}+a_{2}\geq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+3=\lceil q_{1}\rceil+\lceil q_{2}\rceil+1$. This results in an upper quota violation which is not possible by Theorem 4.4 since there exists a lower quota violation by assumption.
If $\sum d_{i}=0$, all of the quotas are integers. However, if the quotas are all integers, this contradicts the assumption that a lower quota violation exists.
If $\sum d_{i}=1,a_{1}+a_{2}\geq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+2\geq\lceil q_{1}\rceil+\lceil q_{2}\rceil$. This is the largest possible value of $a_{1}+a_{2}$ so equality must hold. Then $a_{3}=M-a_{1}-a_{2}=\sum\lfloor q_{i}\rfloor+\sum d_{i}-(\lfloor q_{1}\rfloor+1)-(\lfloor q_{2}\rfloor+1)=\lfloor q_{3}\rfloor-1$. The final apportionment is then $a_{1}=\lceil q_{1}\rceil$, $a_{2}=\lceil q_{2}\rceil$, and $a_{3}=\lfloor q_{3}\rfloor-1$ and, as this is the only possible case, the claim is proven.
In the above lemma, the state with the lower quota violation was arbitrary. The next lemma will justify calling it state three when populations are ordered $p_{1}<p_{2}<p_{3}$:
Lemma 5.3.
In Adams’, Webster’s, and Huntington-Hill’s methods for $n=3$ states, lower quota violations can only occur on the state with the largest population.
Proof.
Begin by ordering the states such that $q_{1}<q_{2}<q_{3}$. Assume for the sake of contradiction there is a lower quota violation on state $j$ for $j=1$ or $2$. Then, by Lemma 5.2 state three is apportioned its upper quota. Thus, the priority value that assigns state three its upper quota is greater than the priority value that would put state $j$ at its lower quota:
$\frac{q_{j}}{d(\lfloor q_{j}\rfloor-1)}<\frac{q_{3}}{d(\lfloor q_{3}\rfloor)}.$
This simplifies to
$q_{j}d(\lfloor q_{3}\rfloor)<q_{3}d(\lfloor q_{j}\rfloor-1)$
Note that $q_{3}>q_{j}>1$ as a lower quota violation is assumed and states are guaranteed at least one seat.
The next lemma, Lemma 5.4, shows the above inequality implies $q_{j}>q_{3}$, a contradiction. ∎
Lemma 5.4.
Let $d$ be the divisor function of Adams’s, Dean’s, or Huntington-Hill’s method. Then for $a,b\in[1,\infty)$, $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor-1)$ implies $a\geq b$.
Proof.
For Dean’s method, let $d(s)=\frac{2s(s+1)}{2s+1}$. First, notice if $a<2$, then $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor)$ never holds. Thus $a\geq 2$. Additionally, this implies the statement is true if $b<2$. The proof for $a,b\geq 2$ will be done by showing the contrapositive. Let $2\leq a\leq b$. Let $n=\lfloor a\rfloor,m=\lfloor b\rfloor$. Note $m\geq n\geq 1$. Define $f(x)=\dfrac{x}{d(\lfloor x\rfloor-1)}$ and $g(x)=\dfrac{x}{d(\lfloor x\rfloor)}$. Then, it suffices to show $f(a)\geq g(b)$.
To do so, first notice for $k\in\mathbb{Z},k\geq 1$, on each interval $[k,k+1)$, both $f(x)$ and $g(x)$ are increasing as their derivatives are positive. Since $f$ is increasing, $f(a)\geq f(n)$ on $[n,n+1)$. Similarly, since $g$ is increasing, $g(b)\leq g(m+1^{-})$ where $m+1^{-}$ indicates approaching $m+1$ from the left on $[m,m+1)$. Then, by simplifying and rearranging:
$g(m+1^{-})=\dfrac{(m+1)(m-0.5)}{m(m-1)}=f(m+1).$
As $m\geq n$, both are integers, and $\dfrac{(m+1)(m-0.5)}{m(m-1)}$ is a decreasing function for $m\in(1,\infty)$. It follows that $f(n)\geq g(m+1^{-})$. Hence, $f(a)\geq f(n)\geq g(m+1^{-})\geq g(b)$.
For the Huntington-Hill method or Adams’s method, the proof is similar.
Summarizing the results of Lemma 5.3 and Lemma 5.2 gives the following:
Theorem 5.5.
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Adams’s, Dean’s, or Huntington-Hill’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
5.2. Lower Quota Violation Criteria Test for 3 States
Next a test is provided to check for lower quota violations on three states. This provides an analytic check for a quota violation which will then be used to find the probability of a lower quota violation for 3 states.
Theorem 5.6.
(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. (1)
    $q_{3}d(\lfloor q_{1}\rfloor)<q_{1}d(\lfloor q_{3}\rfloor-1)$
  2. (2)
    $q_{3}d(\lfloor q_{2}\rfloor)<q_{2}d(\lfloor q_{3}\rfloor-1)$
  3. (3)
    $\lceil q_{1}\rceil+\lceil q_{2}\rceil+\lfloor q_{3}\rfloor=M+1$ (equivalently for $q_{1},q_{2}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1)$.
Proof.
First assume both $q_{1},q_{2}\notin\mathbb{Z}$. Next, assume $\lfloor q_{i}\rfloor>0$ for all $i$. Then the Lower Quota Violation Criteria Test is equivalent to:
  1. (1)
    $\dfrac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\dfrac{q_{1}}{d\lfloor q_{1}\rfloor)}$
  2. (2)
    $\dfrac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\dfrac{q_{2}}{d\lfloor q_{2}\rfloor)}$
  3. (3)
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1$.
To prove this biconditional statement, first assume a lower quota violation from $A(p_{1},p_{2},p_{3})$. Then, by Theorem 5.5, the apportionment is $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$.
Then the priority value to apportion state three its lower quota is less than the priority values to apportion states one and two their upper quota. Thus,
$\frac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\frac{q_{1}}{d(\lfloor q_{1}\rfloor)}\text{ \,and\, }\frac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\frac{q_{2}}{d(\lfloor q_{2}\rfloor)}.$
Finally, from the apportionment,
$M=a_{1}+a_{2}+a_{3}=\lceil q_{1}\rceil+\lceil q_{2}\rceil+\lfloor q_{3}\rfloor-1=\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor+1$
and Condition (3) then follows.
Now assume (1)-(3) are true and continue to assume $\lfloor q_{i}\rfloor>0$. Start to compute $A(p_{1},p_{2},p_{3})$= $A(q_{1},q_{2},q_{3})$ and consider the apportionment algorithm after all states are apportioned $\lfloor q_{i}\rfloor-1$ seats (guaranteed by by Lemma 5.2). Then, by (3), $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1$ and there are 4 seats left to apportion. By (1) and (2), the priority values of states 1 and 2 at their ceilings are higher than the priority value of state 3 at its floor. Thus, the remaining four seats are apportioned to states 1 and 2 and the final apportionment is $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, resulting in a lower quota violation.
If $\lfloor q_{1}\rfloor=0$ (or similarly for state 2), then $\lceil q_{1}\rceil=1$ and the priority for giving state 1 its first seat is infinite. In that case, the corresponding inequality is vacuously true and the same counting argument above applies. The very large priority ensures state 1 receives its ceiling (=1), and the rest of the argument is unchanged.
If $q_{1}$ or $q_{2}$ are integers, then the proof is similar. ∎
5.3. Probability Function for Lower Quota Violations in Three States
For a fixed $M$, $n=3$ states, and with $A$ Adams’s, Dean’s, or the Huntington-Hill method, this section uses the Lower Quota Violation Criteria Test (5.6) to find the probability of a lower quota violation.
Additionally, as the set $\{(q_{1},q_{2})\,|\,q_{1}\text{ or }q_{2}\in\mathbb{Z}\}$ has area zero, assume that quotas of the smallest two states, $q_{1},q_{2}$ are not integers. Therefore, Condition (3) of the Lower Quota Violation Criteria Test (5.6) is then $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1$.
5.3.1. Motivating Example
As an example, let $M=100,\lfloor q_{1}\rfloor=1,\lfloor q_{2}\rfloor=2$, and $A$ the Huntington-Hill method. Write $q_{1}=\lfloor q_{1}\rfloor+d_{1}$ and $q_{2}=\lfloor q_{2}\rfloor+d_{2}$, where $d_{1},d_{2}$ are the decimal parts. The feasible region of the Lower Quota Violation Criterion Test is graphed below;
The probability of a quota violation in the case that $\lfloor q_{1}\rfloor=1$, $\lfloor q_{2}\rfloor=2$, with $q_{1}$ and $q_{2}$ assumed to be uniformly distributed, is the area of the feasible region.
Note the feasible region could be empty as in the example with $M=100,\lfloor q_{1}\rfloor=20,\lfloor q_{2}\rfloor=27$, and $A$ the Huntington-Hill method as shown below:
5.3.2. Calculation of the Feasible Region
The triangular shape of the feasible region in the examples above is not a coincidence. The next theorem shows this is always the case as the feasible region is bounded by three linear inequalities.
Theorem 5.7.
(Equations of the Lower Quota Violation Feasible Region) For $M$ seats, $n=3$ states, and standard quotas $q_{1},q_{2}$, define $R=M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor$. Then the feasible region of the Lower Quota Violation Criteria Test (5.6) is the subset inside the unit square $(0,1)\times(0,1)$ defined by the following linear inequalities:
$\begin{array}[]{clc}y&>\left(\dfrac{d(R-2)+d(\lfloor q_{1}\rfloor)}{-d(\lfloor q_{1}\rfloor)}\right)x+\dfrac{Rd(\lfloor q_{1}\rfloor)-\lfloor q_{1}\rfloor d(R-2)}{d(\lfloor q_{1}\rfloor)}&(1)\\[15.00002pt] y&>\left(\dfrac{-d(\lfloor q_{2}\rfloor)}{d(R-2)+d(\lfloor q_{2}\rfloor)}\right)x+\dfrac{Rd(\lfloor q_{2}\rfloor)-\lfloor q_{2}\rfloor d(R-2)}{d(R-2)+d(\lfloor q_{2}\rfloor)}&(2)\\[15.00002pt] y&<1-x&(3)\end{array}$
with division by 0 interpreted as a vertical line.
Proof.
This is a brute force calculation with given floors $\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor$. Assume $\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor>0$ and $d(\lfloor q_{1}\rfloor),d(\lfloor q_{2}\rfloor)\neq 0$. Write the quotas as:
$(q_{1},q_{2},q_{3})=(\lfloor q_{1}\rfloor+d_{1},\lfloor q_{2}\rfloor+d_{2},M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-d_{1}-d_{2}),$
where $d_{1},d_{2}\in[0,1)$ are the decimal parts of $q_{1},q_{2}$.
Substituting condition (3) of the Lower Quota Violation Criteria Test (5.6) gives
$\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-d_{1}-d_{2}\rfloor=M-1.$
If $d_{1}+d_{2}\leq 1$, then the floor of the third term is $M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-1$ and the equality is satisfied. If $d_{1}+d_{2}>1$, the third term is $M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-2$ violating condition (3). Thus a lower quota violation is possible only if $d_{1}+d_{2}\leq 1$.
Conditions (1) and (2) of the Lower Quota Violation Criteria Test become:
$\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-d_{1}-d_{2}\right)d\left(\lfloor q_{1}\rfloor\right)<\left(\lfloor q_{1}\rfloor+d_{1}\right)d\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-2\right),$
$\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-d_{1}-d_{2}\right)d\left(\lfloor q_{2}\rfloor\right)<\left(\lfloor q_{2}\rfloor+d_{2}\right)d\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-2\right).$
For condition (1), using $R=M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor$, solve for $d_{2}$:
$d_{2}>\left(\frac{d(M-R-2)+d(\lfloor q_{1}\rfloor)}{-d(\lfloor q_{1}\rfloor)}\right)d_{1}+\frac{Rd(\lfloor q_{1}\rfloor)-\lfloor q_{1}\rfloor d(R-2)}{d(\lfloor q_{1}\rfloor)}$
Similarly, solving for $d_{2}$ in condition (2) gives:
$d_{2}>\left(\frac{-d(\lfloor q_{2}\rfloor)}{d(R-2)+d(\lfloor q_{2}\rfloor)}\right)d_{1}+\frac{Rd(\lfloor q_{2}\rfloor)-\lfloor q_{2}\rfloor d(R-2)}{d(R-2)+d(\lfloor q_{2}\rfloor)}$
The feasible region, when non-empty, is the overlap of the above linear inequalities defined over $(0,1)\times(0,1)$, representing all possible decimal parts of the quotas $q_{1},q_{2}$.
If $d(\lfloor q_{1}\rfloor)=0$ or $d(\lfloor q_{2}\rfloor)=0$, then the respective inequality degenerates to a vertical line, but the feasible region remains well defined. ∎
Remark 5.8.
Denote the feasible region defined above as $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$. It satisfies the following properties:
  1. (1)
    (Subset of the quota feasible region) For Dean’s, Adam’s, or Huntington-Hill’s method, and for all pairs ($\lfloor q_{1}\rfloor$$\lfloor q_{2}\rfloor$), $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)\subseteq\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Indeed if a point in $\bigtriangleup_{l}$ lay outside this set, a lower quota violation would occur for a state that does not have the largest population, contradicting Lemma 5.3.
  2. (2)
    (Triangular shape) For $M>3$ the intersection of the three linear inequalities defining $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ is a triangle. This is because the slope of the first line is always less than 1-1, the slope of the second line is always greater than 1-1, and the slope of the third line is always 1-1.
  3. (3)
    (Contained in the Unit Square) If the divisor function $d$ satisfies $\frac{s}{d(s-1)}\geq\frac{s^{\prime}}{d(s^{\prime})}$ for all $s,s^{\prime}\in\mathbb{N}$ and $s^{\prime}\leq s$, then the region bounded by the linear inequalities is a proper subset of the unit square. Geometrically, this is because the $y$-intercept of the first linear inequality is always greater than $1$ and the $x$-intercept of the second linear inequality is always greater than $1$, ensuring the triangular intersection is inside the unit square whenever it is nonempty.
    Algebraically, the $y$-intercept condition
    $\frac{d\left(\lfloor q_{1}\rfloor\right)\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor\right)-\lfloor q_{1}\rfloor d\left(M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor-2\right)}{d\left(\lfloor q_{1}\rfloor\right)}<1$
    rearranges to
    $\frac{\lfloor q_{1}\rfloor}{d(\lfloor q_{1}\rfloor)}>\frac{\lfloor q_{3}\rfloor}{d(\lfloor q_{3}\rfloor-1)},$
    where then $M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor=\lfloor q_{3}\rfloor+1$. By symmetry, the same holds for the $x$-intercept. In particular, this holds in Adam’s method, Dean’s method, and the Huntington-Hill Method.
These properties are used implicitly in Corollary 5.9.1.
5.3.3. Probability Formula for Lower Quota Violations
The probability of a lower quota violation of $A(q_{1},q_{2},q_{3})$ for a fixed $q_{1},q_{2}$ is then the double integral over the feasible region of the joint pdf. Summing this probability for for all possible values of $\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor$ gives the following formula:
Theorem 5.9.
(Probability of Lower Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and for apportionment method $A$ equal to Adams, Dean, or Huntington-Hill, the probability of a lower quota violation is given by:
$\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
where $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ denotes the feasible region of the Lower Quota Violation Criteria Test (5.6) defined in Theorem 5.7.
Corollary 5.9.1.
Let $(q_{1},q_{2})$ be chosen uniformly from $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Then, for Dean’s, Adams’s, and Huntington-Hill’s method with $M>3$, the probability of a quota violation is:
[…]
Where, for $R=M-{\lfloor q_{1}\rfloor}-{\lfloor q_{2}\rfloor}$, $X_{i}$ and $Y_{i}$ are defined by:
$\begin{array}[]{cl}X_{1}&=\dfrac{d\left({\lfloor q_{2}\rfloor}\right)+d\left(R-2\right)-Rd\left({\lfloor q_{2}\rfloor}\right)+{\lfloor q_{2}\rfloor}d\left(R-2\right)}{d\left(R-2\right)}\\[10.00002pt] Y_{1}&=\displaystyle{1-X_{1}}\\[10.00002pt] Y_{2}&=\dfrac{d({\lfloor q_{1}\rfloor})+d(R-2)-Rd({\lfloor q_{1}\rfloor})+{\lfloor q_{1}\rfloor}d(R-2)}{d(R)}\\ X_{2}&=\displaystyle{1-Y_{2}}\\[10.00002pt] X_{3}&=\dfrac{{\lfloor q_{2}\rfloor}d({\lfloor q_{1}\rfloor})+Rd({\lfloor q_{1}\rfloor})-{\lfloor q_{1}\rfloor}d({\lfloor q_{2}\rfloor})-{\lfloor q_{1}\rfloor}d(R-2)}{d(R-2)+d({\lfloor q_{2}\rfloor})+d({\lfloor q_{1}\rfloor})}\\[10.00002pt] Y_{3}&=\dfrac{{\lfloor q_{1}\rfloor}d({\lfloor q_{2}\rfloor})+Rd({\lfloor q_{2}\rfloor})-\lfloor q_{2}\rfloor d({\lfloor q_{1}\rfloor})-{\lfloor q_{2}\rfloor}d(R-2)}{d(R-2)+d({\lfloor q_{2}\rfloor})+d\left({\lfloor q_{1}\rfloor}\right)}.\end{array}$
Additionally, if $M=3$, the probability of a lower quota violation is $\frac{1}{3}$.
Proof.
For each fixed $\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor$, the area of $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ under a uniform measure is $\iint_{\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+x)\,dx\,dy$ . The vertices $(X_{i},Y_{i})$ of $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ are given by the formulas above. The area of $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ is then:
[…]
Where the right hand rule and the location of $(X_{3},Y_{3})$ determines whether the region is nonempty. Since the total area of $\{(q_{1},q_{2})\mid q_{1},q_{2}<M-q_{1}-q_{2}\}$ is $\frac{M^{2}}{6}$, the stated probability formula follows.
For $M=3$, in all three methods
$\mathrm{Area}(\bigtriangleup_{l}(0,1))=\mathrm{Area}(\bigtriangleup_{l}(1,0))=0,\mathrm{Area}(\bigtriangleup_{l}(0,0))=\frac{1}{2}.$
The total probability is then $\frac{1}{2}\cdot\frac{6}{3^{2}}=\frac{1}{3}$. ∎
5.3.4. Empirical Results
Given an apportionment method $A$, $M$ seats, and $n=3$ states, the probability of a lower quota violation under a uniform distribution is computed using Corollary 5.9.1. The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
Quota Violations in the Huntington-Hill Method with Uniform Quotas
$M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
33 $0.33333$ $0.33533$ $\left(0.33240,0.33826\right)$
55 $0.13092$ $0.13055$ $\left(0.12846,0.13264\right)$
$10$ $0.04904$ $0.04915$ $\left(0.04781,0.05049\right)$
$15$ $0.02916$ $0.02906$ $\left(0.02802,0.03010\right)$
2020 $0.02047$ $0.02069$ $\left(0.01981,0.02157\right)$
$100$ $0.00328$ $0.00351$ $\left(0.00314,0.00388\right)$
Quota Violations in Adams’s Method with Uniform Quotas
$M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
33 $0.33333$ $0.33249$ $(0.32957,0.33541)$
55 $0.16000$ $0.15852$ $(0.15626,0.16078)$
$10$ $0.11057$ $0.1099$ $(0.10796,0.11184)$
$15$ $0.09897$ $0.09961$ $(0.09775,0.10147)$
2020 $0.09365$ $0.09143$ $(0.08964,0.09322)$
$100$ $0.08210$ $0.08188$ $(0.08018,0.08358)$
Quota Violations in Dean’s Method with Uniform Quotas
$M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
33 $0.33333$ $0.33363$ $(0.33071,0.33655)$
55 $0.13714$ $0.13818$ $(0.13604,0.14032)$
$10$ $0.05581$ $0.05545$ $(0.05403,0.05687)$
$15$ $0.03397$ $0.03412$ $(0.03299,0.03525)$
2020 $0.02403$ $0.02444$ $(0.02348,0.02540)$
$100$ $0.00376$ $0.00362$ $(0.00325,0.00399)$
The code used for random sampling is available at https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git.
6. Structure and Probability of Upper Quota Violations in Jefferson’s Method for Three States
6.1. Analogous Theoretical Results
We now provide a similar study of upper quota violations in Jefferson’s method when apportioning three states. Webster’s, Adams’s, Dean’s, and the Huntington-Hill Method have no upper quota violations by Lemma 5.1 so only Jefferson’s method needs to be considered.
Lemma 6.1.
In a divisor apportionment method with $M$ seats, $n=3$ states, and standard quotas $q_{i}$, if an upper quota violation occurs then the apportionment of the state with the violation is $\lceil q_{i}\rceil+1$, and the other states are assigned their lower quotas.
Proof.
Assume without loss of generality that the upper quota violation occurs on state 3. The claim is that the final apportionment is then $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$. Since the first two states cannot be apportioned less than the floor of the quota by Theorem 4.4, $a_{1}\geq\lfloor q_{1}\rfloor$ and $a_{2}\geq\lfloor q_{2}\rfloor$. Additionally, as an upper quota violations is assume $a_{2}\geq\lceil q_{3}\rceil+1$.
Write $q_{i}=\lfloor q_{i}\rfloor+d_{i}$, where $d_{i}\in[0,1)$ is the decimal part. Then, in particular $\sum d_{i}\in\{0,1,2\}$.
Then, assume $d_{3}\neq 0$. As such, $\lfloor q_{3}\rfloor=\lceil q_{3}\rceil-1$ and
$a_{1}+a_{2}=M-a_{3}\leq M-\lceil q_{i}\rceil-1=\sum\lfloor q_{i}\rfloor+\sum d_{i}-\lceil q_{i}\rceil-1\leq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\sum d_{i}-2.$
Then, for each case of $\sum d_{i}\in\{0,1,2\}$:
If $\sum d_{i}=0$, then all of the quotas are integers and the apportionment is not a quota violation.
If $\sum d_{i}=1$, then $a_{1}+a_{2}\leq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor-1$, implying that one of states one or two is assigned less than their lower quota, which is a contradiction with the earlier statement $a_{1}\geq\lfloor q_{1}\rfloor$ and $a_{2}\geq\lfloor q_{2}\rfloor$.
If $\sum d_{i}=2$, then $a_{1}+a_{2}\leq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor$. As it is also the case that $a_{1}\geq\lfloor q_{1}\rfloor$ and $a_{2}\geq\lfloor q_{2}\rfloor$, this suggests that they are all in fact equal. Then, $a_{3}=M-a_{1}-a_{2}=\sum\lfloor q_{i}\rfloor+\sum d_{i}-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor=\lfloor q_{3}\rfloor+2=\lceil q_{3}\rceil+1$. Thus, $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$ and as this is the only possible case for $d_{3}\neq 0$ the claim is proven under this assumption.
Then, if $d_{3}=0$:
$a_{1}+a_{2}=M-a_{3}\leq M-\lceil q_{i}\rceil-1=\sum\lfloor q_{i}\rfloor+\sum d_{i}-\lceil q_{i}\rceil-1\leq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\sum d_{i}-1.$
Which gives for each case of $\sum d_{i}\in\{0,1,2\}$:
If $\sum d_{i}=0$, then all of the quotas are integers and the apportionment is not a quota violation.
If $\sum d_{i}=2$, then as $d_{i}\in[0,1)$ this violates the assumption that $d_{3}=0$.
If $\sum d_{i}=1$, then $a_{1}+a_{2}\leq\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor$. As it is also the case that $a_{1}\geq\lfloor q_{1}\rfloor$ and $a_{2}\geq\lfloor q_{2}\rfloor$, this suggests that they are all in fact equal. Then, $a_{3}=M-a_{1}-a_{2}=\sum\lfloor q_{i}\rfloor+\sum d_{i}-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor=\lfloor q_{3}\rfloor+1=\lceil q_{3}\rceil+1$. Thus, $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$ and as this is the only possible case for $d_{3}=0$ the claim is proven here as well. ∎
Next, similar to Lemma 5.3, we justify using $q_{3}$ in the above proof as having the upper quota violation.
Lemma 6.2.
In Jefferson’s method and $n=3$ states, upper quota violations can only occur on the state with the largest population.
Proof.
Order the state’s populations $p_{1}<p_{2}<p_{3}$. Assume there is an upper quota violation on state $i$ with $i\in\{1,2\}$. Lemma 6.1 then implies both that state three is not apportioned $\lceil q_{3}\rceil$ as state $i$ is apportioned $\lceil q_{i}\rceil+1.$ This gives the following relation between priority values:
$\frac{q_{i}}{d(\lceil q_{i}\rceil)}>\frac{q_{3}}{d(\lfloor q_{3}\rfloor)}$
Rearranging gives:
$q_{i}(\lfloor q_{3}\rfloor+1)>q_{3}(\lceil q_{i}\rceil+1)$
The next Lemma will show that this implies $q_{i}>q_{3}$, a contradiction. ∎
Lemma 6.3.
For $a,b\in(0,\infty)$, $a(\lfloor b\rfloor+1)>b(\lceil a\rceil+1)$ implies $a>b$.
Proof.
If $b\in\mathbb{Z}$, then the statement is $a(b+1)>b(\lceil a\rceil+1)$, which rearranges to $ab+a>b\lceil a\rceil+b\geq ab+b$, hence $a>b$. If $b$ is not an integer, the statement rearranges to $a\lceil b\rceil>b\lceil a\rceil+b\geq\lfloor b\rfloor a+b$. Then, $b<a(\lceil b\rceil-\lfloor b\rfloor)=a$. ∎
Summarizing Lemma 6.1 and 6.2 gives the following:
Theorem 6.4.
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
6.1.1. Upper Quota Violation Criteria Test for 3 States
Next a test is provided to check for an upper quota violation on three states under Jefferson Apportionment. Similar to Theorem 5.6, this provides an analytic and more direct way to check for a quota violation without needing to run the apportionment algorithm. This test will then be used to find the probability of a upper quota violation for 3 states.
Theorem 6.5.
(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. (1)
    $q_{3}d(\lfloor q_{1}\rfloor)>q_{1}d(\lceil q_{3}\rceil)$
  2. (2)
    $q_{3}d(\lfloor q_{2}\rfloor)>q_{2}d(\lceil q_{3}\rceil)$
  3. (3)
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
The proof is similar to the proof of Theorem 5.6.
Proof.
Assume that there is an upper quota violation. Then, by Theorem 6.4, the apportionment is $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$. Then, states one and two are no assigned the floors of their quotas and state three is assigned the ceiling of its quota plus one and writing the relationship between priority values suggested by this grants conditions one and two. For condition three, $M=a_{1}+a_{2}+a_{3}=\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil+1$.
Assume that conditions one through three hold. Then, calculate the apportionment by first assigning every state $\lfloor q_{i}\rfloor$, where every state is guaranteed at least that many seats by Theorem 4.5. Then, assign the remaining seats by following the priority value relations implied by conditions one and two, which is assigning all additional seats to state three. This creates the apportionment $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, which is an upper quota violation. ∎
From there, in analogy with Theorem 5.7, derive the following probability result:
Theorem 6.6.
(Equations of the Upper Quota Violation Feasible Region) For $M$ seats, $n=3$ states, and standard quotas $q_{1},q_{2}$, define $R=M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor$. Then the feasible region of the Upper Quota Violation Criteria Test (6.5) is the subset inside the unit square $(0,1)\times(0,1)$ defined by the following linear inequalities:
$\begin{array}[]{clc}\displaystyle{y}&<\displaystyle{\left(\dfrac{d(R-1)+d(\lfloor q_{1}\rfloor)}{-d(\lfloor q_{1}\rfloor)}\right)x+\dfrac{Rd(\lfloor q_{1}\rfloor)-\lfloor q_{1}\rfloor d(R-1)}{d\left(\lfloor q_{1}\rfloor\right)}}&(1)\\[15.00002pt] y&<\displaystyle{\left(\frac{-d(\lfloor q_{2}\rfloor)}{d(R-1)+d(\lfloor q_{2}\rfloor)}\right)x+\frac{Rd(\lfloor q_{2}\rfloor)-\lfloor q_{2}\rfloor d(R-1)}{d(R-1)+d(\lfloor q_{2}\rfloor)}}&(2)\\[15.00002pt] y&>\displaystyle{1-x}&(3)\end{array}$
where division by 0 is interpreted as a vertical line.
The proof is omitted as it is similar to the proof of Theorem 5.7. Similar properties of Remark 5.8 apply to this region.
6.1.2. Probability Formula for an Upper Quota Violation
For a fixed number of seats $M$, the probability of an upper quota violation of $A(q_{1},q_{2},q_{3})$ for a fixed $q_{1},q_{2}$ is then the double integral over the feasible region of the joint pdf. Summing this probability for for all possible values of $\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor$ gives the following formula:
Theorem 6.7.
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
$\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
where $\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ denotes the feasible region of the Upper Quota Violation Criteria Test (6.5) defined in Theorem 6.6.
Corollary 6.7.1.
If $(q_{1},q_{2})$ are chosen uniformly on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$, then for Jefferson’s method the probability of an quota violation for any $M$ is given by:
[…]
Where $X_{i}$ and $Y_{i}$ are defined, with $R=M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor$, as:
$\begin{array}[]{cl}X_{1}&=\dfrac{Rd({\lfloor q_{1}\rfloor})-{\lfloor q_{1}\rfloor}d(R-1)-d({\lfloor q_{1}\rfloor})}{d(R-1)}\\[10.00002pt] Y_{1}&=\displaystyle{1-X_{1}}\\[10.00002pt] Y_{2}&=\dfrac{Rd({\lfloor q_{2}\rfloor})-{\lfloor q_{2}\rfloor}d(R-1)-d({\lfloor q_{2}\rfloor})}{d(R-1)}\\[10.00002pt] X_{2}&=\displaystyle{1-Y_{2}}\\[10.00002pt] X_{3}&=\dfrac{Rd({\lfloor q_{1}\rfloor})-{\lfloor q_{1}\rfloor}d(R-1)-{\lfloor q_{1}\rfloor}d({\lfloor q_{2}\rfloor})+{\lfloor q_{2}\rfloor}d({\lfloor q_{1}\rfloor})}{d(R-1)+d({\lfloor q_{1}\rfloor})+d({\lfloor q_{2}\rfloor})}\\[10.00002pt] Y_{3}&=\dfrac{Rd({\lfloor q_{2}\rfloor})-{\lfloor q_{2}\rfloor}d(R-1)-{\lfloor q_{2}\rfloor}d({\lfloor q_{1}\rfloor})+{\lfloor q_{1}\rfloor}d({\lfloor q_{2}\rfloor})}{d(R-1)+d({\lfloor q_{1}\rfloor})+d({\lfloor q_{2}\rfloor})}.\end{array}$
The proof is similar to that of Corollary 5.9.1.
6.1.3. Empirical Results
Given Jefferson’s method, $M$ seats, and $n=3$ states, the probability of an upper quota violation under a uniform distribution is computed using Corollary 6.7.1. The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
Quota Violations in Jefferson’s Method with Uniform Quotas
$M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
33 $0.02222$ $0.02279$ $(0.02187,0.02371)$
55 $0.03943$ $0.03888$ $(0.03768,0.04008$
$10$ $0.05671$ $0.05694$ $(0.05550,0.05838)$
$15$ $0.06362$ $0.06114$ $(0.05966,0.06262)$
2020 $0.06729$ $0.06693$ $(0.06538,0.06848)$
$100$ $0.07686$ $0.07548$ $(0.07384,0.07712)$
The code used for random sampling is available at https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git.
7. Conclusion and Next Steps
This work provides a comprehensive probabilistic analysis of quota violations in divisor methods for three states, advancing the quantitative understanding of the Balinski–Young Theorem. By classifying the structure of both lower and upper quota violations, deriving analytic criteria tests, and developing exact probability formulas, we obtain an understanding how often these apportionment violations occur and under what conditions. The combination of theoretical results with empirical validation provides a foundation for understanding the nature of violations where both the number of states and seats are increased.
The results are significant because they illustrate how probabilistic and geometric techniques are combined to analyze apportionment. Specifically, these methods both quantify the frequency and illuminate the structure of quota violations in divisor methods that are foundational in political science and economics. The ability to move beyond impossibility theorems toward concrete probability calculations represents a step forward in the applied probability analysis of voting and apportionment rules.
Several natural directions for future research emerge from this study. First, the methods developed here invite generalization to larger numbers of states and all divisor methods, where the geometry of quota violations becomes richer and potentially more intricate. Second, extending the framework beyond divisor methods to rank-index methods or even to the entire class of apportionment methods covered by the Balinski–Young Theorem would provide a unified probabilistic perspective on fairness violations in apportionment. Third, exploring alternative population distributions, including those motivated by empirical census data, would connect the theoretical findings more closely to the applied settings. Finally, a computational analysis where the number of seats or states grows large could offer insights into the limiting behavior of quota violation probabilities.
The results presented here make a meaningful contribution to the theory and practice of apportionment. They not only clarify the mathematical structure of quota violations in divisor methods but also open new avenues for probabilistic, geometric, and computational investigations. We believe that the framework developed here will serve as a foundation for further research at the intersection of applied probability, social choice theory, and political representation.
8. Declarations
8.1. Ethical Approval and Consent to Participate
Not applicable.
8.2. Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8.3. Data Availability Statement
The raw data generated during this study is available upon request. All relevant processed data is included within this article and the appendix. The Python code is available to run here: https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git
8.4. Consent for publication
Not applicable.
8.5. Funding
This research was supported by Johns Hopkins’ Provost’s Undergraduate Research Award (PURA)

### v3

@@
 Quantifying the Balinski-Young Theorem: Structure and Probability of Quota Violations in Divisor Methods for Three States
-Abstract.
+Abstract
 The apportionment problem asks how to assign representation to states based on their populations.
 That is, given census data and a fixed number of seats, how many seats should each state be assigned?
-Various algorithms exist to solve the apportionment problem, but by the Balinski-Young Theorem, every such algorithm will be flawed in some way.
+Various algorithms exist to solve the apportionment problem, but by the Balinski-Young Impossibility Theorem, every such algorithm will be flawed in some way.
 This paper focuses on divisor methods of apportionment, where the possible flaws are known as quota violations.
-This paper presents a detailed analysis of quota violations that can arise under divisor methods for three states.
-The study focuses on quota violations in Adams’s, Jefferson’s, Dean’s, and the Huntington-Hill methods when allocating $M$ seats.
-Theoretical results are proved about the behavior of these methods, particularly focusing on the types of quota violations that may occur, their frequency, and their structure.
-The paper then introduces tests to detect quota violations, which are then employed to construct a probability function which calculates the likelihood of such violations occurring given an initial three state population vector whose components follow varying distributions.
-1. Introduction
+This paper presents a detailed analysis of quota violations that can arise under divisor methods for three states, By focusing on the three-state case, the paper makes the consequences of the Balinski–Young theorem particularly transparent and allows for a precise classification of quota violations that is difficult to obtain in more general formulations.
+The study focuses on quota violations in the Adams, Jefferson, Dean, and the Huntington-Hill methods when allocating $M$ seats, but is expandable to a wider class of divisor functions.
+Theoretical results are proved about the behavior of these methods, particularly focusing on the types of quota violations that may occur, their frequency, and their structure and geometry.
+The key results of the paper are tests to detect quota violations which are employed to understand the geometry of violations and construct a probability function which calculates the likelihood of such violations occurring given an initial three state population vector whose components follow varying distributions.
+1 Introduction
 The apportionment problem - dividing a fixed number of indivisible resources among states in proportion to their populations - lies at the intersection of political science, combinatorics, and probability.
-A key class of apportionment methods are divisor methods, which are a group of simple methods that avoid unfair results known as apportionment paradoxes [BY01].
-Divisor methods have been widely employed in the United States for assigning seats in House of Representatives, with Jefferson’s, Webster’s method, or the Huntington-Hill methods used for most of the country’s history.
+A key class of apportionment methods are divisor methods, which are a group of simple methods that avoid unfair results known as apportionment paradoxes [undefa].
+Divisor methods have been widely employed in the United States for assigning seats in House of Representatives, with the Jefferson, Webster, or Huntington-Hill method used for most of the country’s history.
 The Huntington-Hill method is currently used to apportion the House of Representatives.
+[undefe]
 While divisor methods avoid paradoxes, these methods do not always satisfy the quota rule, which requires that each state receive an allocation equal to either the floor or ceiling of its exact proportional share.
-This phenomenon of imperfect apportionment is in fact guaranteed by the Balinski-Young Theorem [BY01], which establishes that no apportionment method can simultaneously satisfy the quota rule and avoid paradoxes.
-Since divisor methods can violate quota, and as flawed methods are unavoidable, a good question to then ask is how often will a quota violation happen, and can quota violations be predicted in advance given states’ populations?
-This paper answers that question for three states.
-Specifically, this paper investigates the frequency and structure of quota violations under divisor methods in three states in the methods of Adams’s, Jefferson’s, Webster’s, Dean’s, and Huntington-Hill.
+This phenomenon of imperfect apportionment is in fact guaranteed by the Balinski-Young Impossibility Theorem [undefa], which establishes that no apportionment method can simultaneously satisfy the quota rule and avoid paradoxes.
+Since divisor methods can violate quota, and as flawed methods are unavoidable, good questions to ask are: What is the underlying structure of such violations?
+How often will a quota violation happen?
+And, can quota violations be predicted in advance given states’ populations?
+This paper answers these question for three states.
+Specifically, The paper investigates the frequency, structure, and geometry of quota violations under divisor methods in three states in the Adams, Jefferson, Webster, Dean, and Huntington-Hill methods.
 The paper characterizes how quota violations occur and provides formulas to calculate the probability that a divisor method violates quota.
-2. Preliminary Notation and Definitions
+Overall, this paper provides a detailed analysis of quota violations in the simpler case of three states, which ultimately highlights and corroborates the Balinski Young Impossibility Theorem.
+This paper builds on related work about probabilities and analysis of apportionment within social choice theory and the mathematics of elections, such as similar probabilistic work on simplexes concerning the Jagiellonian compromise found in [undefo] and [undefn]; work on systemic bias such as in [undefl], [undefm], [undefb], and [undefh]; and work on the probabilities of voting paradoxes such as in [undefc] and [undefj].
+2 Preliminary Notation and Definitions
 Definition 2.1.
-Given a strictly increasing divisor function $d:\mathbb{N}\to\mathbb{R}_{\geq 0}$, populations $(p_{1},\dots,p_{n})$ and a fixed number of seats $M$, an apportionment function $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$, is then defined as follows:
+(Divisor Method) Given a strictly increasing divisor function $d:\mathbb{Z}_{\geq 0}\to\mathbb{R}_{\geq 0}$, populations $(p_{1},\dots,p_{n})$ and a fixed number of seats $M$, an apportionment function $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$, is then defined as follows:
   1. 1.
@@
 Equivalently, if $d(0)=0$, replace the first step with assigning each state one seat rather than zero.
-The most common and arguably reasonable divisor methods are the “five workable methods” [BY77, BY01, Hunt21] defined by the following divisor functions:
+The most common and arguably reasonable divisor methods are the “five workable methods” [undef, undefa, undefg] defined by the following divisor functions:
 Name Divisor Function
@@
 Informally, a quota violation is when a state is apportioned too many or too few seats.
-The Balinski-Young Theorem [BY01] guarantees that all apportionment methods either have quota violations or have strange, unfair occurrences known as apportionment paradoxes.
-Divisor methods have quota violations but avouid paradoxes [BY01].
-3. Results
-The main results of the paper are:
+The Balinski-Young Theorem [undefa] guarantees that all apportionment methods either have quota violations or have strange, unfair occurrences known as apportionment paradoxes.
+Divisor methods have quota violations but avoid paradoxes [undefa].
+3 Results
+The following is a summary of the results of the paper:
 Theorem 5.1.
 For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur.
-In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
+In particular, upper quota violations do not occur in the Adams, Dean, and Huntington-Hill methods for $n=3$ states.
 Theorem 5.5.
 (Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$.
-If $A$ is Adams’s, Dean’s, or the Huntington Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$.
+If $A$ is the Adams, Dean, or the Huntington Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$.
 That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
 Theorem 5.6.
-(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$.
+(Lower Quota Violation Criteria Test) Let $A$ be the Adams, Dean, or the Huntington-Hill method, with divisor function $d(s)$.
 Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$.
 Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
-  1. (1)
+  1. 1.
     $q_{3}d(\lfloor q_{1}\rfloor)<q_{1}d(\lfloor q_{3}\rfloor-1)$
-  2. (2)
+  2. 2.
     $q_{3}d(\lfloor q_{2}\rfloor)<q_{2}d(\lfloor q_{3}\rfloor-1)$
-  3. (3)
+  3. 3.
     $\lceil q_{1}\rceil+\lceil q_{2}\rceil+\lfloor q_{3}\rfloor=M+1$ (equivalently for $q_{1},q_{2}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1)$.
@@
 (Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$.
-If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$.
+If $A$ is the Jefferson method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$.
 That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
 Theorem 6.5.
-(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$.
+(Upper Quota Violation Criteria Test) Let $A$ be the Jefferson method, with divisor function $d(s)$.
 Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$.
 Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
-  1. (1)
+  1. 1.
     $q_{3}d(\lfloor q_{1}\rfloor)>q_{1}d(\lceil q_{3}\rceil)$
-  2. (2)
+  2. 2.
     $q_{3}d(\lfloor q_{2}\rfloor)>q_{2}d(\lceil q_{3}\rceil)$
-  3. (3)
-    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
+  3. 3.
+    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{3}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
 Theorem 6.7.
 (Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$.
-For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
+For $q_{1},q_{2}<q_{3}$ and the Jefferson method, the probability of a upper quota violation is given by:
 $\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
 where $\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ denotes the feasible region of the Upper Quota Violation Criteria Test (6.5) defined in Theorem 6.6.
-4. Background Information
+4 Background Information
 The following are simple lemmas, known results, and common assumptions that are utilized in this paper.
@@
 Let $A(p_{1},\ldots,p_{n})$ be the divisor method apportionment for state populations $p_{i}$, with $A$ an arbitrary divisor method.
-4.1. Assumptions
+4.1 Assumptions
 Assume there are no ties in priority values; in particular this includes assuming no ties in states’ populations. Additionally, to guarantee at least one seat per state assume $M\geq n$.
-4.2. Introductory Lemmas and Definitions
+4.2 Introductory Lemmas and Definitions
 Lemma 4.1.
@@
 Lemma 4.2.
-(Homogeneity) A divisor method is preserved under proportionality.
+[undefa] (Homogeneity) A divisor method is preserved under proportionality.
 If $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$ then for any scalar, $\lambda\in\mathbb{R}_{>0}$, $A(\lambda p_{1},\dots,\lambda p_{n})=(a_{1},\dots,a_{n})$.
 In particular, apportioning the standard quotas is the same as apportioning the populations: $A(p_{1},\dots,p_{n})=A(q_{1},\dots,q_{n})$.
-[BY01].
+.
 Definition/Lemma 4.3.
 Define a divisor apportionment method $\tilde{A}(p_{1},\ldots,p_{n})$ on $n$ states and $M$ seats by finding a modified divisor $D\in\mathbb{R}_{>0}$ such that $D$ causes $\sum_{i=0}^{n}a_{i}=M$, for $a_{i}$ calculated as follows:
-  1. (1)
+  1. 1.
     For each state, calculate a modified quota $m_{i}=\frac{p_{i}}{D}.$
-  2. (2)
+  2. 2.
     For each state, assign seats by modified-rounding: round up if $m_{i}\geq d(\lfloor m_{i}\rfloor)$; round down if less.
-The equivalence of the apportionment methods $A(p_{1},\ldots,p_{n})$ in Definition 2.1 and $\tilde{A}(p_{1},\ldots,p_{n})$ is classical. See [BY01], [Hunt21], and [Pukelsheim2017].
+The equivalence of the apportionment methods $A(p_{1},\ldots,p_{n})$ in Definition 2.1 and $\tilde{A}(p_{1},\ldots,p_{n})$ is classical. See [undefa], [undefg], and [undefk].
 Theorem 4.4.
-No divisor method will have both upper and lower quota violations for any $M$. [BY01, KS18].
+[undefa, undefi] No divisor method will have both upper and lower quota violations for any $M$.
 Theorem 4.5.
-(No Violation Theorem) From [BY01], a summary of cases when no quota violations occur is as follows:
+[undefa] (No Violation Theorem) From , a summary of cases when no quota violations occur is as follows:
   • (a)
@@
   • (b)
-    For $n=3$ states, Webster’s Method is the unique method such that upper or lower quota violations do not occur.
+    For $n=3$ states, the Webster method is the unique method such that upper or lower quota violations do not occur.
   • (c)
-    Adams’s Method is the unique method where upper quota violations do not occur for all $n$.
+    The Adams Method is the unique method where upper quota violations do not occur for all $n$.
   • (d)
-    Jefferson’s Method is the unique method where lower quota violations do not occur for all $n$.
-5. Structure and Probability of Quota Violations for Three States in Adams’s, Dean’s, and Huntington-Hill’s Method
-We start by providing a detailed study of lower quota violations in Adams’s, Dean’s, and Huntington-Hill’s method when apportioning three states.
-Webster’s and Jefferson’s method have no lower quota violations by Theorem 4.5.
-5.1. Characterization of Quota Violations in Three States
-By the following theorem, only lower quota violations can occur in Adams’s, Dean’s, and Huntington-Hill’s method for three states.
+    The Jefferson Method is the unique method where lower quota violations do not occur for all $n$.
+5 Structure and Probability of Quota Violations for Three States in the Adams, Dean, and Huntington-Hill Methods
+We start by providing a detailed study of lower quota violations in the Adams, Dean, and Huntington-Hill methods when apportioning three states.
+The Webster and Jefferson methods have no lower quota violations by Theorem 4.5.
+5.1 Characterization of Quota Violations in Three States
+By the following theorem, only lower quota violations can occur in the Adams, Dean, and Huntington-Hill methods for three states.
 Theorem 5.1.
 For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur.
-In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
+In particular, upper quota violations do not occur in the Adams, Dean, and Huntington-Hill methods for $n=3$ states.
 Proof.
@@
 The proof above allows for lower quota violations to occur in $n=3$ states and indeed they do.
-The following three lemmas completely classify the lower quota violations in Adams’s, Dean’s and Huntington-Hill’s apportionment methods.
+The following three lemmas completely classify the lower quota violations in the Adams, Dean and Huntington-Hill apportionment methods.
 Lemma 5.2.
@@
 Lemma 5.3.
-In Adams’, Webster’s, and Huntington-Hill’s methods for $n=3$ states, lower quota violations can only occur on the state with the largest population.
+In the Adams, Webster, and Huntington-Hill methods for $n=3$ states, lower quota violations can only occur on the state with the largest population.
 Proof.
@@
 Lemma 5.4.
-Let $d$ be the divisor function of Adams’s, Dean’s, or Huntington-Hill’s method. Then for $a,b\in[1,\infty)$, $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor-1)$ implies $a\geq b$.
+Let $d$ be the divisor function of the Adams, Dean, or Huntington-Hill method. Then for $a,b\in[1,\infty)$, $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor-1)$ implies $a\geq b$.
 Proof.
-For Dean’s method, let $d(s)=\frac{2s(s+1)}{2s+1}$.
+For the Dean method, let $d(s)=\frac{2s(s+1)}{2s+1}$.
 First, notice if $a<2$, then $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor)$ never holds.
@@
 Hence, $f(a)\geq f(n)\geq g(m+1^{-})\geq g(b)$.
-For the Huntington-Hill method or Adams’s method, the proof is similar.
+For the Huntington-Hill method or the Adams method, the proof is similar.
 Summarizing the results of Lemma 5.3 and Lemma 5.2 gives the following:
@@
 (Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$.
-If $A$ is Adams’s, Dean’s, or Huntington-Hill’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$.
+If $A$ is the Adams, Dean, or Huntington-Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$.
 That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
-5.2. Lower Quota Violation Criteria Test for 3 States
+5.2 Lower Quota Violation Criteria Test for 3 States
 Next a test is provided to check for lower quota violations on three states.
@@
 Theorem 5.6.
-(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$.
+(Lower Quota Violation Criteria Test) Let $A$ be the Adams, Dean, or Huntington-Hill method, with divisor function $d(s)$.
 Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$.
 Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
-  1. (1)
+  1. 1.
     $q_{3}d(\lfloor q_{1}\rfloor)<q_{1}d(\lfloor q_{3}\rfloor-1)$
-  2. (2)
+  2. 2.
     $q_{3}d(\lfloor q_{2}\rfloor)<q_{2}d(\lfloor q_{3}\rfloor-1)$
-  3. (3)
+  3. 3.
     $\lceil q_{1}\rceil+\lceil q_{2}\rceil+\lfloor q_{3}\rfloor=M+1$ (equivalently for $q_{1},q_{2}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1)$.
@@
 First assume both $q_{1},q_{2}\notin\mathbb{Z}$. Next, assume $\lfloor q_{i}\rfloor>0$ for all $i$. Then the Lower Quota Violation Criteria Test is equivalent to:
-  1. (1)
+  1. 1.
     $\dfrac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\dfrac{q_{1}}{d\lfloor q_{1}\rfloor)}$
-  2. (2)
+  2. 2.
     $\dfrac{q_{3}}{d(\lfloor q_{3}\rfloor-1)}<\dfrac{q_{2}}{d\lfloor q_{2}\rfloor)}$
-  3. (3)
+  3. 3.
     $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1$.
@@
 If $q_{1}$ or $q_{2}$ are integers, then the proof is similar. ∎
-5.3. Probability Function for Lower Quota Violations in Three States
-For a fixed $M$, $n=3$ states, and with $A$ Adams’s, Dean’s, or the Huntington-Hill method, this section uses the Lower Quota Violation Criteria Test (5.6) to find the probability of a lower quota violation.
+Remark 5.7 (Generalization to $n$ states).
+Although Theorem 5.6 is stated for three states, the underlying ideas extend in principle to $n\geq 4$.
+The main complication already appears at $n=4$: there is no analogue of Theorem 5.5 providing a single canonical form for quota violations.
+Instead, one must account for multiple distinct configurations, each requiring its own criterion, leading to a disjunction of conditions.
+The number of such cases grows rapidly with $n$, making the resulting criterion unwieldy.
+For clarity and focus, we restrict attention to the case $n=3$ and leave the general $n$-state setting for future work.
+5.3 Geometric Visualization of Lower Quota Violations
+Using the Lower Quota Violation Criteria Test (5.6) and Theorem 5.8, we can display all quota violations on the simplex of standard quotas $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$ within the Adams, Dean, or Huntington-Hill method.
+This creates a visualization of the geometry and structure of lower quota violations.
+For instance, below is a simplex plot of all violatory $(q_{1},q_{2},q_{3})$ in the Huntington-Hill method with $M=13$:
+Figure 1: Quota Violations in the Huntington-Hill Method on $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$
+5.4 Probability Function for Lower Quota Violations in Three States
+For a fixed $M$, $n=3$ states, and with $A$ the Adams, Dean, or Huntington-Hill method, this section uses the Lower Quota Violation Criteria Test (5.6) to find the probability of a lower quota violation.
 Additionally, as the set $\{(q_{1},q_{2})\,|\,q_{1}\text{ or }q_{2}\in\mathbb{Z}\}$ has area zero, assume that quotas of the smallest two states, $q_{1},q_{2}$ are not integers.
 Therefore, Condition (3) of the Lower Quota Violation Criteria Test (5.6) is then $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-1$.
-5.3.1. Motivating Example
+5.4.1 Motivating Example
 As an example, let $M=100,\lfloor q_{1}\rfloor=1,\lfloor q_{2}\rfloor=2$, and $A$ the Huntington-Hill method.
 Write $q_{1}=\lfloor q_{1}\rfloor+d_{1}$ and $q_{2}=\lfloor q_{2}\rfloor+d_{2}$, where $d_{1},d_{2}$ are the decimal parts.
-The feasible region of the Lower Quota Violation Criterion Test is graphed below;
+The feasible region of the Lower Quota Violation Criterion Test is graphed below:
+Figure 2: Example Inequality Region with Overlap
 The probability of a quota violation in the case that $\lfloor q_{1}\rfloor=1$, $\lfloor q_{2}\rfloor=2$, with $q_{1}$ and $q_{2}$ assumed to be uniformly distributed, is the area of the feasible region.
 Note the feasible region could be empty as in the example with $M=100,\lfloor q_{1}\rfloor=20,\lfloor q_{2}\rfloor=27$, and $A$ the Huntington-Hill method as shown below:
-5.3.2. Calculation of the Feasible Region
+Figure 3: Example Inequality Region without Overlap
+5.4.2 Calculation of the Feasible Region
 The triangular shape of the feasible region in the examples above is not a coincidence. The next theorem shows this is always the case as the feasible region is bounded by three linear inequalities.
@@
 Denote the feasible region defined above as $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$. It satisfies the following properties:
-  1. (1)
-    (Subset of the quota feasible region) For Dean’s, Adam’s, or Huntington-Hill’s method, and for all pairs ($\lfloor q_{1}\rfloor$$\lfloor q_{2}\rfloor$), $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)\subseteq\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$.
+  1. 1.
+    (Subset of the quota feasible region) For the Dean, Adam, or Huntington-Hill methods, and for all pairs ($\lfloor q_{1}\rfloor$$\lfloor q_{2}\rfloor$), $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)\subseteq\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$.
 Indeed if a point in $\bigtriangleup_{l}$ lay outside this set, a lower quota violation would occur for a state that does not have the largest population, contradicting Lemma 5.3.
-  2. (2)
+  2. 2.
     (Triangular shape) For $M>3$ the intersection of the three linear inequalities defining $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)$ is a triangle.
 This is because the slope of the first line is always less than 1-1, the slope of the second line is always greater than 1-1, and the slope of the third line is always 1-1.
-  3. (3)
+  3. 3.
     (Contained in the Unit Square) If the divisor function $d$ satisfies $\frac{s}{d(s-1)}\geq\frac{s^{\prime}}{d(s^{\prime})}$ for all $s,s^{\prime}\in\mathbb{N}$ and $s^{\prime}\leq s$, then the region bounded by the linear inequalities is a proper subset of the unit square.
@@
 By symmetry, the same holds for the $x$-intercept.
-In particular, this holds in Adam’s method, Dean’s method, and the Huntington-Hill Method.
+In particular, this holds in Adam’s method, Dean method, and the Huntington-Hill Method.
 These properties are used implicitly in Corollary 5.10.1.
-5.3.3. Probability Formula for Lower Quota Violations
+5.4.3 Probability Formula for Lower Quota Violations
 The probability of a lower quota violation of $A(q_{1},q_{2},q_{3})$ for a fixed $q_{1},q_{2}$ is then the double integral over the feasible region of the joint pdf.
@@
 Let $(q_{1},q_{2})$ be chosen uniformly from $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$.
-Then, for Dean’s, Adams’s, and Huntington-Hill’s method with $M>3$, the probability of a quota violation is:
+Then, for the Dean, Adams, and Huntington-Hill method with $M>3$, the probability of a quota violation is:
 […]
@@
 The total probability is then $\frac{1}{2}\cdot\frac{6}{3^{2}}=\frac{1}{3}$. ∎
-5.3.4. Empirical Results
+5.4.4 Empirical Results
 Given an apportionment method $A$, $M$ seats, and $n=3$ states, the probability of a lower quota violation under a uniform distribution is computed using Corollary 5.10.1.
-The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
+The resulting probabilities are shown in the tables below and compared with a simulation of 100,000 random $(q_{1},q_{2})$ samples from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
 Quota Violations in the Huntington-Hill Method with Uniform Quotas
@@
 $100$ $0.00328$ $0.00351$ $\left(0.00314,0.00388\right)$
-Quota Violations in Adams’s Method with Uniform Quotas
+Quota Violations in the Adams Method with Uniform Quotas
 $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
@@
 $100$ $0.08210$ $0.08188$ $(0.08018,0.08358)$
-Quota Violations in Dean’s Method with Uniform Quotas
+Quota Violations in the Dean Method with Uniform Quotas
 $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
@@
 The code used for random sampling is available at https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git.
-6. Structure and Probability of Upper Quota Violations in Jefferson’s Method for Three States
-6.1. Analogous Theoretical Results
-We now provide a similar study of upper quota violations in Jefferson’s method when apportioning three states.
-Webster’s, Adams’s, Dean’s, and the Huntington-Hill Method have no upper quota violations by Lemma 5.1 so only Jefferson’s method needs to be considered.
+Comparisons to simulations under different distributions may be found in Appendix A.
+6 Structure and Probability of Upper Quota Violations in the Jefferson Method for Three States
+6.1 Analogous Theoretical Results
+We now provide a similar study of upper quota violations in the Jefferson method when apportioning three states.
+The Webster, Adams, Dean, and the Huntington-Hill methods have no upper quota violations by Lemma 5.1 so only the Jefferson method needs to be considered.
 Lemma 6.1.
@@
 Lemma 6.2.
-In Jefferson’s method and $n=3$ states, upper quota violations can only occur on the state with the largest population.
+In the Jefferson method and $n=3$ states, upper quota violations can only occur on the state with the largest population.
 Proof.
@@
 (Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$.
-If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$.
+If $A$ is the Jefferson method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$.
 That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
-6.1.1. Upper Quota Violation Criteria Test for 3 States
+6.1.1 Upper Quota Violation Criteria Test for 3 States
 Next a test is provided to check for an upper quota violation on three states under Jefferson Apportionment.
@@
 Theorem 6.5.
-(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$.
+(Upper Quota Violation Criteria Test) Let $A$ be the Jefferson method, with divisor function $d(s)$.
 Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$.
 Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
-  1. (1)
+  1. 1.
     $q_{3}d(\lfloor q_{1}\rfloor)>q_{1}d(\lceil q_{3}\rceil)$
-  2. (2)
+  2. 2.
     $q_{3}d(\lfloor q_{2}\rfloor)>q_{2}d(\lceil q_{3}\rceil)$
-  3. (3)
-    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
+  3. 3.
+    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{3}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
 The proof is similar to the proof of Theorem 5.6.
@@
 ∎
-From there, in analogy with Theorem 5.7, derive the following probability result:
+From there, in analogy with Theorem 5.8, derive the following result:
 Theorem 6.6.
@@
 where division by 0 is interpreted as a vertical line.
-The proof is omitted as it is similar to the proof of Theorem 5.7. Similar properties of Remark 5.8 apply to this region.
-6.1.2. Probability Formula for an Upper Quota Violation
+The proof is omitted as it is similar to the proof of Theorem 5.8. Similar properties to Remark 5.9 apply to this region.
+6.1.2 Geometric Visualization of Upper Quota Violations
+Similarly to Section 5.3, one may use Theorem 6.5 and specifically Theorem 6.6 to display all violatory $(q_{1},q_{2},q_{3})$ on the simplex $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$.
+This allows for a visualization of the geometry and structure of upper quota violations.
+For instance, below is a simplex plot indicating all violatory $(q_{1},q_{2},q_{3})$ in the Jefferson method with $M=13$:
+Figure 4: Quota Violations in the Jefferson Method on $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$
+6.1.3 Probability Formula for Upper Quota Violations
 For a fixed number of seats $M$, the probability of an upper quota violation of $A(q_{1},q_{2},q_{3})$ for a fixed $q_{1},q_{2}$ is then the double integral over the feasible region of the joint pdf.
@@
 (Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$.
-For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
+For $q_{1},q_{2}<q_{3}$ and the Jefferson method, the probability of a upper quota violation is given by:
 $\sum_{\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor}\iint_{\bigtriangleup_{u}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)}f(\lfloor q_{1}\rfloor+x,\lfloor q_{2}\rfloor+y)\,dx\,dy$
@@
 Corollary 6.7.1.
-If $(q_{1},q_{2})$ are chosen uniformly on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$, then for Jefferson’s method the probability of an quota violation for any $M$ is given by:
+If $(q_{1},q_{2})$ are chosen uniformly on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$, then for the Jefferson method the probability of an quota violation for any $M$ is given by:
 […]
@@
 The proof is similar to that of Corollary 5.10.1.
-6.1.3. Empirical Results
-Given Jefferson’s method, $M$ seats, and $n=3$ states, the probability of an upper quota violation under a uniform distribution is computed using Corollary 6.7.1.
+6.1.4 Empirical Results
+Given the Jefferson method, $M$ seats, and $n=3$ states, the probability of an upper quota violation under a uniform distribution is computed using Corollary 6.7.1.
 The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
-Quota Violations in Jefferson’s Method with Uniform Quotas
+Quota Violations in the Jefferson Method with Uniform Quotas
 $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
@@
 The code used for random sampling is available at https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git.
-7. Conclusion and Next Steps
-This work provides a comprehensive probabilistic analysis of quota violations in divisor methods for three states, advancing the quantitative understanding of the Balinski–Young Theorem.
-By classifying the structure of both lower and upper quota violations, deriving analytic criteria tests, and developing exact probability formulas, we obtain an understanding how often these apportionment violations occur and under what conditions.
-The combination of theoretical results with empirical validation provides a foundation for understanding the nature of violations where both the number of states and seats are increased.
+Comparisons to simulations under different distributions may be found in Appendix A.
+7 Conclusion and Next Steps
+This work provides a comprehensive analysis of quota violations in divisor methods for three states, advancing the quantitative and qualitative understanding of the Balinski–Young Theorem.
+The main contribution of the paper lies in making the consequences of the Balinski–Young theorem explicit and quantitative in this restricted setting by providing a complete classification of possible quota violations and analyzing the conditions and frequencies under which they occur.
+The work thus refines and illustrates a classical impossibility result rather than challenging it or proposing an alternative apportionment method
 The results are significant because they illustrate how probabilistic and geometric techniques are combined to analyze apportionment.
-Specifically, these methods both quantify the frequency and illuminate the structure of quota violations in divisor methods that are foundational in political science and economics.
+Specifically, these methods both quantify the frequency and illuminate the structure of quota violations in divisor methods.
 The ability to move beyond impossibility theorems toward concrete probability calculations represents a step forward in the applied probability analysis of voting and apportionment rules.
 Several natural directions for future research emerge from this study.
-First, the methods developed here invite generalization to larger numbers of states and all divisor methods, where the geometry of quota violations becomes richer and potentially more intricate.
+First and foremost, a strong and in progress next step of this work is the generalization to larger numbers of states and all divisor methods, where the geometry of quota violations becomes richer and more intricate.
 Second, extending the framework beyond divisor methods to rank-index methods or even to the entire class of apportionment methods covered by the Balinski–Young Theorem would provide a unified probabilistic perspective on fairness violations in apportionment.
-Third, exploring alternative population distributions, including those motivated by empirical census data, would connect the theoretical findings more closely to the applied settings.
-Finally, a computational analysis where the number of seats or states grows large could offer insights into the limiting behavior of quota violation probabilities.
+Finally, a computational analysis where the number of seats, states or ratios between populations grows large could offer insights into the limiting behavior of quota violation probabilities.
 The results presented here make a meaningful contribution to the theory and practice of apportionment.
@@
 We believe that the framework developed here will serve as a foundation for further research at the intersection of applied probability, social choice theory, and political representation.
-8. Declarations
-8.1. Ethical Approval and Consent to Participate
+8 Declarations
+8.1 Ethical Approval and Consent to Participate
 Not applicable.
-8.2. Competing Interests
+8.2 Competing Interests
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-8.3. Data Availability Statement
+8.3 Data Availability Statement
 The raw data generated during this study is available upon request.
@@
 The Python code is available to run here: https://github.com/TylerCWunder/Probability_and_Structure_of_Quota_Violations_in_Divisor_Methods_of_Apportionment_code.git
-8.4. Consent for publication
+8.4 Consent for publication
 Not applicable.
-8.5. Funding
+8.5 Funding
 This research was supported by Johns Hopkins’ Provost’s Undergraduate Research Award (PURA)
+Appendix A Comparison of Theoretical Results to Simulations Under Various Distributions
+The following are the results of additional simulations and comparisons to theoretical values calculated using Theorems 5.10 and 6.7.
+Probability of Quota Violations with $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(\frac{1}{2},\frac{1}{2},\frac{1}{2})$
+Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
+Huntington-Hill 33 $0.55051$ $0.5527$ $(0.54962,0.55578)$
+Adams 33 $0.55051$ $0.55421$ $(0.55113,0.55729)$
+Dean 33 $0.55051$ $0.54921$ $(0.54613,0.55229)$
+Jefferson 33 $0.01252$ $0.01283$ $(0.01213,0.01353)$
+Huntington-Hill 55 $0.33758$ $0.33468$ $(0.33176,0.33760)$
+Adams 55 $0.37385$ $0.37233$ $(0.36933,0.37533)$
+Dean 55 $0.34624$ $0.34727$ $(0.34432,0.35022)$
+Jefferson 55 $0.03003$ $0.02934$ $(0.02829,0.03039)$
+Huntington-Hill $10$ $0.20313$ $0.20478$ $(0.20228,0.20728)$
+Adams $10$ $0.29367$ $0.29473$ $(0.29190,0.29756)$
+Dean $10$ $0.21604$ $0.21577$ $(0.21322,0.21831)$
+Jefferson $10$ $0.05637$ $0.05682$ $(0.05539,0.05825)$
+Huntington-Hill $15$ $0.15545$ $0.15436$ $(0.15212,0.15660)$
+Adams $15$ $0.26701$ $0.26744$ $(0.26470,0.27018)$
+Dean $15$ $0.16703$ $0.1651$ $(0.16280,0.16740)$
+Jefferson $15$ $0.07108$ $0.07055$ $(0.06896,0.07214)$
+Probability of Quota Violations with $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(2,2,2)$
+Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
+Huntington-Hill 33 $0.13580$ $0.13519$ $(0.13307,0.13731)$
+Adams 33 $0.13580$ $0.13486$ $(0.13274,0.13698)$
+Dean 33 $0.13580$ $0.13587$ $(0.13375,0.13799)$
+Jefferson 33 $0.02670$ $0.02627$ $(0.02528,0.02726)$
+Huntington-Hill 55 $0.02276$ $0.2327$ $(0.02234,0.02420)$
+Adams 55 $0.03660$ $0.03729$ $(0.03612,0.03846)$
+Dean 55 $0.02513$ $0.02497$ $(0.02400,0.02594)$
+Jefferson 55 $0.02889$ $0.02862$ $(0.02759,0.02965)$
+Huntington-Hill $10$ $0.00386$ $0.00391$ $(0.00352,0.00430)$
+Adams $10$ $0.03091$ $0.0304$ $(0.02934,0.03146)$
+Dean $10$ $0.00537$ $0.0052$ $(0.00475,0.00565)$
+Jefferson $10$ $0.03045$ $0.03104$ $(0.02997,0.03211)$
+Huntington-Hill $15$ $0.00150$ $0.00154$ $(0.00130,0.00178)$
+Adams $15$ $0.03107$ $0.03195$ $(0.03086,0.03304)$
+Dean $15$ $0.00227$ $0.00242$ $(0.00212,0.00272)$
+Jefferson $15$ $0.03094$ $0.03015$ $(0.02909,0.03121)$
+Remark A.1.
+As a Dirichlet $(1,1,1)$ distributions is a uniform distribution on the simplex $\{(x_{1},x_{2},x_{3})\mid x_{1}+x_{2}+x_{3}=1\}$, the case $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(1,1,1)$ is equivalent to the case presented in Sections 5.4.4 and 6.1.4.
+Probability of Quota Violations with $p_{i}$ IID and $p_{i}\sim\exp(1)$
+Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
+Huntington-Hill 33 $0.33333$ $0.33413$ $(0.33121,0.33705)$
+Adams 33 $0.33333$ $0.33298$ $(0.33006,0.33590)$
+Dean 33 $0.33333$ $0.33379$ $(0.33087,0.33671)$
+Jefferson 33 $0.02222$ $0.02199$ $(0.02108,0.02290)$
+Huntington-Hill 55 $0.13092$ $0.13225$ $(0.13015,0.13434)$
+Adams 55 $0.16000$ $0.15939$ $(0.15712,0.16166)$
+Dean 55 $0.13714$ $0.13545$ $(0.13333,0.13757)$
+Jefferson 55 $0.03943$ $0.03942$ $(0.03821,0.04063)$
+Huntington-Hill $10$ $0.04904$ $0.04949$ $(0.04815,0.05083)$
+Adams $10$ $0.11057$ $0.11116$ $(0.10921,0.11311)$
+Dean $10$ $0.05581$ $0.05663$ $(0.05520,0.05806)$
+Jefferson $10$ $0.05671$ $0.05642$ $(0.05499,0.05785)$
+Remark A.2.
+The probabilities in the case $p_{i}$ are independent and identically distributed random variables with $p_{i}\sim\exp(\lambda)$ are the same as the probabilities presented in sections 5.4.4 and 6.1.4.
+This may be shown by computing the associated probability density function for $(q_{1},q_{2})$.
+Probability of Quota Violations with $p_{i}$ IID and $p_{i}\sim U(0,1000)$
+Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
+Huntington-Hill 33 $0.125$ $0.12422$ $(0.122218,0.12626)$
+Adams 33 $0.125$ $0.12399$ $(0.12195,0.12603)$
+Dean 33 $0.125$ $0.1251$ $(0.12305,0.12715)$
+Jefferson 33 $0.01389$ $0.01341$ $(0.01270,0.1412)$
+Huntington-Hill 55 $0.03923$ $0.03966$ $(0.03845,0.04087)$
+Adams 55 $0.05903$ $0.05944$ $(0.05797,0.06091)$
+Dean 55 $0.04360$ $0.04444$ $(0.04316,0.04572)$
+Jefferson 55 $0.01819$ $0.01802$ $(0.01720,0.01884)$
+Huntington-Hill $10$ $0.01363$ $0.01357$ $(0.01285,0.01429)$
+Adams $10$ $0.04015$ $0.03966$ $(0.03845,0.04087)$
+Dean $10$ $0.01612$ $0.01571$ $(0.01494,0.01648)$
+Jefferson $10$ $0.02308$ $0.02278$ $(0.02186,0.02370)$
+Huntington-Hill $15$ $0.00802$ $0.00807$ $(0.00752,0.00862)$
+Adams $15$ $0.03664$ $0.03724$ $(0.03607,0.03841)$
+Dean $15$ $0.00953$ $0.01006$ $(0.00944,0.01068)$
+Jefferson $15$ $0.02536$ $0.02561$ $(0.02463,0.02659)$